OpenStudy (anonymous):

3+(4x - 5)= 2/3(6x-3)

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

x = ?

OpenStudy (anonymous):

x = 0.375

OpenStudy (texaschic101):

check the problem again Alexander....its not 3 times..it is 3 plus

OpenStudy (alexandervonhumboldt2):

-2+4x=4x-2

OpenStudy (alexandervonhumboldt2):

0=0

OpenStudy (anonymous):

0.0

OpenStudy (texaschic101):

3 + (4x - 5) = 2/3(6x - 3) = 4x - 2 = 4x - 2 0 = 0 infinite solutions

OpenStudy (anonymous):

so the answer is 0?

OpenStudy (anonymous):

no the answer is infinite